op lxk)*(kxi) is: The result ofplease show work and explain i dont understand at all
Answer #1

The vector product of two vectors is given by

  \vec{A} \times \vec{B}= \begin{vmatrix} \hat{i} & \hat{j} & \hat{k} \\ A_{x} & A_{y} & A_{z} \\ B_{x} & B_{y} & B_{z} \end{vmatrix}

Where

\vec{A}= A_{x}\, \hat{i}+A_{y}\, \hat{j}+A_{z}\,\hat{k}

and   \vec{B}= B_{x}\, \hat{i}+B_{y}\, \hat{j}+B_{z}\,\hat{k}

Now, the unit vectors \hat{i} , \hat{j} and \hat{k} can be writtten as

  \hat{i} = 1. \hat{i}+0. \hat{j} +0. \hat{k}

\hat{j} = 0. \hat{i}+1. \hat{j} +0. \hat{k}  

and \hat{k} = 0. \hat{i}+0. \hat{j} +1. \hat{k}

So,

  \hat{j} \times \hat{k}= \begin{vmatrix} \hat{i} & \hat{j} & \hat{k} \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{vmatrix}= \hat{i} .(1-0)- \haat{j}.(0-0)+\hat{k}.(0-0)= \hat{i}

Similarly,

\hat{k} \times \hat{i}= \begin{vmatrix} \hat{i} & \hat{j} & \hat{k} \\ 0 & 0 & 1 \\ 1 & 0 & 0 \end{vmatrix}= \hat{i} .(0-0)- \haat{j}.(0-1)+\hat{k}.(0-0)= \hat{j}

And,

  \hat{i} \times \hat{j}= \begin{vmatrix} \hat{i} & \hat{j} & \hat{k} \\ 1 & 0 & 0 \\ 0 & 1 & 0 \end{vmatrix}= \hat{i} .(0-0)- \haat{j}.(0-0)+\hat{k}.(1-0)= \hat{k}

Therefore,

  (\hat{j} \times \hat{k}) \times (\hat{k} \times \hat{i})= \hat{k}